If someone was to run for 46minutes at a speed of 9 kilometers per hour, how far did he run in miles?

Answer by mananth(16949)

46 minutes = 46/60 hours = 23/30 hours
9km/hr = 5.59 mph ( 1km/h =0.62 mph)
D= rate * time
5.59*23/30
=4.28 miles